OpenStudy (anonymous):

Tony Stark's personal assistant Pepper tries to escape an automated drone by running horizontally with a speed of 6.00 m.s off the roof of a 35.0 m tall building. Tony Stark who has been standing right below the edge of the roof, tries to catch her about 1.5 m above street level. Ignore air resistance. How fast must he be running at a constant speed in order to catch pepper? How far from the building will he catch her?

OpenStudy (anonymous):

The time it will take Pepper to fall (35-1.5)m= 33.5m = 0.5 g t^2 = 0.5 (9.8) t^2. t = sqrt(6.84)= 2.6 s. Stark needs to run along at the same horizontal velocity that Pepper had, 6.0m/s. He will be x=v t=(6.0)(2.6)= 15.6m from the building. She will hit with a bang!